# Use of Nash equilibrium in finding game theoretic robust security bound on quantum bit error rate

### v1

@@ Abstract
Nash equilibrium is employed to find a game theoretic robust security bound on quantum bit error rate (QBER) for DL04 protocol which is a scheme for quantum secure direct communication that has been experimentally realized recently. The receiver, sender, and eavesdropper (Eve) are considered to be quantum players (players having the capability to perform quantum operations). Specifically, Eve is considered to have the capability of performing quantum attacks (e.g., Wójcik’s original attack, Wójcik’s symmetrized attack, and Pavičić attack) and classical intercept and resend attack. Game theoretic analysis of the security of DL04 protocol in the above scenario is performed by considering several sub-game scenarios. The analysis revealed the absence of a Pareto optimal Nash equilibrium point within these sub-games. Consequently, mixed strategy Nash equilibrium points are identified and employed to establish both upper and lower bounds for QBER. Further, the vulnerability of the DL04 protocol to Pavičić attack in the message mode is established. In addition, it is observed that the quantum attacks performed by Eve are more powerful than the classical attack, as the QBER value and the probability of detecting Eve’s presence are found to be lower in quantum attacks compared to classical ones.
@@ I Introduction
Game theory examines and models how individuals behave in situations involving strategic (rational) thinking and interactive decision-making. It is crucial for decision-making processes and assessing opportunities, both in business and everyday scenarios. Instances requiring strategic thinking are prevalent in fields such as economics [1], political science [2], biology [3, 4], and military applications [5, 6]. Participants in these scenarios have their own sets of potential actions, referred to as strategies, and express preferences for these actions through a payoff matrix. Game theory is concerned with representing these activities and identifying optimal strategies. Among the various concepts in game theory, the Nash equilibrium is particularly significant. It characterizes the optimal decisions considering the actions of other players. In a Nash equilibrium, no player stands to benefit by altering their strategy alone [7, 8].
@@
Quantum mechanics stands out as one of the most influential theories throughout history. Despite the controversies it has sparked since its inception, its predictions have been consistently and precisely confirmed through experiments [9]. Quantum game theory enables the examination of interactive decision-making scenarios involving players utilizing quantum technology. This technology serves a dual purpose: functioning as a quantum communication protocol and providing a more efficient method for randomizing players’ strategies compared to classical games [10]. Quantum game theory emanated in 1999 through the contributions of David Meyer [11] and Jens Eisert, Martin Wilkens, and Maciej Lewenstein [12]. Their research explored games that incorporated quantum information, showcasing scenarios where quantum players demonstrated advantages over their classical counterparts. Subsequently, numerous examples of quantum games, largely built upon the foundations established by Meyer and Eisert, Wilkens, and Lewenstein, have been extensively examined. Further details and references can be found in surveys such as [13]. In the realm of experiments, researchers have successfully implemented the quantum version of the Prisoners’ Dilemma game using an NMR quantum computer [14]. Additionally, Vaidman [15] demonstrated a simple game wherein players consistently emerge victorious if they share a GHZ state beforehand, contrasting with classical players where winning is always subject to probability. Quantum strategies have been employed to introduce fairness elements into remote gambling scenarios [16] and in formulating algorithms for quantum auctions, which come with numerous security advantages [17]. Flitney and Abbott [18] have explored quantum adaptations of Parrondo’s games. Such analyses not only aid in designing secure networks leading to the identification of novel quantum algorithms but also add an entirely distinct dimension to characterizing a game or a protocol [19, 20]. Furthermore, eavesdropping [21, 22] and optimal cloning [23] can be conceptualized as games played between participants.
@@ II Preliminaries of Game Theory
Before we delve into the technical details of our work, it will be apt to briefly introduce the basics of game theory. Game theory comprises a set of mathematical models designed to examine scenarios involving both competition and collaboration, where an individual’s (player’s) ability to make choices effectively relies on the choices made by others. Each player has a set of possible strategies or actions they can take. This strategy is a plan or decision that specifies how a player will act in different situations within the game. The primary objective is to identify optimal strategies for individuals facing such situations and to identify equilibrium states. The outcomes of a game are often represented in terms of payoffs, which measure the utility or satisfaction that each player receives based on the chosen strategies of all players. Payoffs can be represented in various forms, such as numerical values, rankings, or other measures. Game theory can be represented in different forms; the normal form (strategic form) is a matrix that shows the payoffs for each combination of strategies chosen by the players. The extensive form (dynamic game) uses a tree-like diagram to represent sequential and simultaneous decision-making. There are mainly two types of games, zero-sum games and non-zero-sum games. In a zero-sum game, the total payoff is constant, and gains for one player result in losses for the other player(s). In non-zero-sum games, the total payoff can vary, and the interests of the players may not be directly opposed. An important fundamental concept in game theory is Nash equilibrium which corresponds to a set of strategies in which no player has an incentive to unilaterally change their strategy, given the strategies chosen by the other players. It represents a stable solution where no player can improve their payoff by changing their strategy. A game’s strategy set is considered Pareto efficient (or Pareto optimal) when there does not exist another strategy set that can improve the outcome for one player without negatively affecting any other player. Further, a dominant strategy is a strategy that is always the best choice for a player, regardless of the strategies chosen by other players. A dominant strategy is a strong concept in rational decision-making. When a player follows a pure strategy, they choose a single action or decision without any randomness or uncertainty. In some games, players may adopt mixed strategies, where they choose their actions with certain probabilities. Mixed strategies can lead to a Nash equilibrium when no pure strategy is optimal [12, 48, 49, 50, 51].
Quantum game In a traditional game that allows for the use of mixed strategies, players construct their strategies by using real coefficients to form convex linear combinations of their pure strategies [52]. In contrast, in a quantum game [12, 53], players employ unitary transformations and quantum states that belong to substantially larger strategy spaces. This has led to discussions proposing that quantum games could be considered extensions of classical games, potentially offering stakeholders a quantum advantage in the game [54]. As mentioned earlier, it is important to emphasize that any quantum protocol (quantum cryptographic scenarios) can be analogized to a game-like scenario. In this context, participants’ strategies hinge on the utilization of unitary operations and the selection of measurement bases, with the final payoff determined by the results of the measurements. Quantum advantage arises from the optimal sequential utilization of quantum operations on quantum states by the players. Typically, in traditional quantum game theory paper classical players are limited to using coherent permutations of standard basis states, or similarly restricted types of unitary operations. In contrast, quantum players have access to a wider range of available unitary operations, which may include the full spectrum of such operations with fewer limitations.
@@
Our contribution In quantum communication protocols, less restricted quantum players can utilize their mixed strategies to calculate their individual best response functions based on their payoffs. The solutions of best response functions give the Nash equilibrium points. While not all of these points may be Pareto optimal Nash equilibrium points, they serve as a basis for determining the probabilities of mixed strategies employed by the players. This, in turn, allows for an investigation into the existence of Pareto optimal Nash equilibrium points. The outcome of this analysis helps to determine the secure bounds for various quantum information parameters (e.g., secret key rate, QBER) with appropriate use of the fundamental concepts of quantum information theory [55, 56, 57, 58]. In our letter, we take the DL04 protocol [24] and analyze it as a game, considering different types of attack (collective attacks and IR attack) to get the secure threshold limit for QBER.
@@
In the DL04 protocol, two quantum players employ strategies to ensure secure communication. Our focus lies in determining the secure bound on QBER when accounting for the presence of an eavesdropper (Eve), who may utilize both quantum and classical attack strategies. We consider that the quantum Eve employs the following attack strategies: Wójcik’s original attack [66], Wójcik’s symmetrized attack [66], and Pavičić attack [67], denoted as $E_{1}$, $E_{2}$, and $E_{3}$, respectively. Additionally, there is a classical Eve employing an Intercept Resend (IR) attack strategy ($E_{4}$). The detailed analysis of attack strategies is conducted in Appendices A, B, C, and D. We will now define the payoff for each party involved. The mutual information between Alice and Bob has a positive impact on the payoffs of legitimate players (Alice and Bob), but it negatively affects Eavesdropper’s (Eve) payoff. Conversely, the mutual information shared between Alice and Eve, as well as Bob and Eve, negatively affects the payoffs of legitimate players. Additionally, legitimate players benefit from the ability to detect Eve’s presence [^1] The contribution of both parameters is equally significant, contingent upon the specific quantitative values of QBER and the detection probability of Eve’s presence ($P_{d}$) observed in both message and control modes of the DL04 protocol., which improves their payoffs in this competitive scenario. On the other hand, Eve’s payoff increases with the acquisition of more information from Alice and Bob but decreases as the mutual information between Alice and Bob increases. Furthermore, Eve incurs penalties if she is detected. Consequently, Eve’s payoff is enhanced when the probability of remaining undetected is higher. Additionally, Eve may employ various quantum gates to collect information from Alice and Bob, and the greater the number of gates she uses, the more it impacts her overhead, which, in turn, adversely affects her payoff. As a result, the payoff can be customized to reflect the various scenarios and benefits of the players that we intend to analyze. The DL04 protocol, resembling a game, can be conceptualized as a zero-sum game. We may formulate the payoffs of Alice, Bob, and Eve for a general attack strategy $\mathcal{E}$ as follows [^2] We assume that the difficulty of preparing quantum states and performing measurement operations is the same for Alice, Bob, and Eve. Therefore, we do not factor in this contribution when estimating the payoffs for these parties.,
@@
where $\omega_{a},\,\omega_{b},\,\omega_{c},\,\omega_{d},\,\omega_{e},\,\omega_{f},\,\omega_{g},\,\omega_{h},\,\omega_{i},\,\omega_{j}$, and $\omega_{k}$ are all positive real numbers, and they are interpreted as the factors or weights assigned to each component within the payoff such that $\stackrel{{\scriptstyle[}}{{m}}=a]{d}{\sum}\omega_{m}=1$ and $\stackrel{{\scriptstyle[}}{{n}}=e]{k}{\sum}\omega_{n}=1$ where $m\in\{a,b,c,d\}$ and $n\in\{e,f,\ldots,k\}$ Further, $I\left({\rm A,B}\right)$, $I\left({\rm A,E}\right)$, and $I\left({\rm B},{\rm E}\right)$ are the mutual information between Alice and Bob, Alice and Eve, and Bob and Eve, respectively. Here, $P_{d}$ is the probability of detection of Eve’s presence, $n_{1}$, $n_{2}$, and $n_{3}$ are the number of single-qubit, two-qubit, and three-qubit gates, respectively. Here, it may be noted that in the payoff function described above, $P_{d}$ and $QBER$ are considered on equal footing because both $P_{d}$ and $QBER$ reveal the presence of Eve with a similar effect.
Considering Eve has unlimited quantum resources and the quantum gates used by her are imperfection-free, we can take $\omega_{i}=\omega_{j}=\omega_{k}=0$ as the above consideration would imply that she can use as many quantum gates as she wants to execute an attack without causing any negative effect on her payoff function. To simplify the analysis, in what follows we have considered the effect of each component in the payoff is the same. Thus, in our consideration, $\omega_{a}=\omega_{b}=\omega_{c}=\omega_{d}=\omega_{e}=\omega_{f}=\omega_{g}=\omega_{h}=0.25$. The Eq. (1) is modified as,

### v2

@@ Abstract
Nash equilibrium is employed to find a game theoretic robust security bound on quantum bit error rate (QBER) for DL04 protocol which is a scheme for quantum secure direct communication that has been experimentally realized recently. The receiver, sender and eavesdropper (Eve) are considered to be quantum players (players having the capability to perform quantum operations). Specifically, Eve is considered to have the capability of performing quantum attacks (e.g., Wójcik’s original attack, Wójcik’s symmetrized attack and Pavičić attack) and classical intercept and resend attack. Game theoretic analysis of the security of DL04 protocol in the above scenario is performed by considering several game scenarios. The analysis revealed the absence of a Pareto optimal Nash equilibrium point within these game scenarios. Consequently, mixed strategy Nash equilibrium points are identified and employed to establish both upper and lower bounds for QBER. Further, the vulnerability of the DL04 protocol to Pavičić attack in the message mode is established. In addition, it is observed that the quantum attacks performed by Eve are more powerful than the classical attack, as the QBER value and the probability of detecting Eve’s presence are found to be lower in quantum attacks compared to classical ones.
@@ I Introduction
Game theory examines and models how individuals behave in situations involving strategic (rational) thinking and interactive decision-making. It is crucial for decision-making processes and assessing opportunities, both in business and everyday scenarios. Instances requiring strategic thinking are prevalent in fields such as economics [1], political science [2], biology [3, 4] and military applications [5, 6]. Participants in these scenarios have their own sets of potential actions, referred to as strategies, and express preferences for these actions through a payoff matrix. Game theory is concerned with representing these activities and identifying optimal strategies. Among the various concepts in game theory, the Nash equilibrium is particularly significant. It characterizes the optimal decisions considering the actions of other players. In a Nash equilibrium, no player stands to benefit by altering their strategy alone [7, 8].
@@
Quantum mechanics stands out as one of the most influential theories throughout history. Despite the controversies it has sparked since its inception, its predictions have been consistently and precisely confirmed through experiments [9]. Quantum game theory enables the examination of interactive decision-making scenarios involving players utilizing quantum technology. This technology serves a dual purpose: functioning as a quantum communication protocol and providing a more efficient method for randomizing players’ strategies compared to classical games [10]. Quantum game theory emanated in 1999 through the contributions of David Meyer [11], and Jens Eisert, Martin Wilkens and Maciej Lewenstein [12]. Their research explored games that incorporated quantum information, showcasing scenarios where quantum players demonstrated advantages over their classical counterparts. Subsequently, numerous examples of quantum games, largely built upon the foundations established by Meyer, and Eisert, Wilkens and Lewenstein, have been extensively examined. Further details and references can be found in surveys such as [13]. In the realm of experiments, researchers have successfully implemented the quantum version of the Prisoners’ Dilemma game using an NMR quantum computer [14]. Additionally, Vaidman [15] demonstrated a simple game wherein players consistently emerge victorious if they share a GHZ state beforehand, contrasting with classical players where winning is always subject to probability. Quantum strategies have been employed to introduce fairness elements into remote gambling scenarios [16] and in formulating algorithms for quantum auctions, which come with numerous security advantages [17]. Flitney and Abbott [18] have explored quantum adaptations of Parrondo’s games. Such analyses not only aid in designing secure networks leading to the identification of novel quantum algorithms but also add an entirely distinct dimension to characterizing a game or a protocol [19, 20, 21]. Furthermore, eavesdropping [22, 23] and optimal cloning [24] can be conceptualized as games played between participants. Thus, an interconnection between game theory and quantum mechanics is already explored. There are primarily two approaches that illustrate this interconnection. In the first approach, quantum resources are used to play a traditional game that can also be played without any quantum resources, but the use of quantum resources provide some advantages, whereas in the second approach, a quantum mechanical scenario is described using the concepts of the game theory. In what follows, the first (second) approach mentioned above is referred to as the quantized game(gaming the quantum). Here, it will be apt to note that a quantized game can be defined as a unitary function mapping the Cartesian product of sets of quantum superpositions of players’ pure strategies to the entire Hilbert space encompassing the domain of the game. This retains the characteristics of the original game under specific conditions. On the other hand, the concept of gaming the quantum [25] refers to the application(s) of the principles of game theory to quantum mechanics to derive game-theoretic solutions. In this study, we do “gaming the quantum” by applying non-cooperative game theory to a quantum communication protocol to demonstrate how Nash equilibrium can serve as a viable solution concept.
@@ II Preliminaries of Game Theory
Before we delve into the technical details of our work, it will be apt to briefly introduce the basics of game theory. Game theory comprises a set of mathematical models designed to examine scenarios involving both competition and collaboration, where an individual’s (player’s) ability to make choices effectively relies on the choices made by others. Each player has a set of possible strategies or actions they can take. This strategy is a plan or decision that specifies how a player will act in different situations within the game. The primary objective is to identify optimal strategies for individuals facing such situations and to identify equilibrium states. The outcomes of a game are often represented in terms of payoffs, which measure the utility or satisfaction that each player receives based on the chosen strategies of all players. Payoffs can be represented in various forms, such as numerical values, rankings or other measures. Game theory can be represented in different forms; the normal form (strategic form) is a matrix that shows the payoffs for each combination of strategies chosen by the players. The extensive form (dynamic game) uses a tree-like diagram to represent sequential and simultaneous decision-making. There are mainly two types of games, zero-sum games and non-zero-sum games. In a zero-sum game, the total payoff is constant, and gains for one player result in losses for the other player(s). In non-zero-sum games, the total payoff can vary, and the interests of the players may not be directly opposed. An important fundamental concept in game theory is Nash equilibrium which corresponds to a set of strategies in which no player has an incentive to unilaterally change their strategy, given the strategies chosen by the other players. It represents a stable solution where no player can improve their payoff by changing their strategy. A game’s strategy set is considered Pareto efficient (or Pareto optimal) when there does not exist another strategy set that can improve the outcome for one player without negatively affecting any other player. Further, a dominant strategy is a strategy that is always the best choice for a player, regardless of the strategies chosen by other players. A dominant strategy is a strong concept in rational decision-making. When a player follows a pure strategy, they choose a single action or decision without any randomness or uncertainty. In some games, players may adopt mixed strategies, where they choose their actions with certain probabilities. Mixed strategies can lead to a Nash equilibrium when no pure strategy is optimal [12, 50, 51, 52, 53].
[…]
Our considered game, which is a non-cooperative quantum game involving multiple players, lacks a pure strategy Nash equilibrium. This is due to the nature of the multiplayer model and the way payoffs are calculated, similar to the approach in [25]. Furthermore, the concept of a Nash equilibrium solution is fundamental in game theory as it can be used to predict the behavior of non-cooperating players. The proof of Nash’s theorem for the existence of an equilibrium in mixed strategies in traditional games is relatively simple and relies entirely on Kakutani’s fixed-point theorem [57]. In the realm of quantum games, Meyer demonstrated the existence of Nash equilibrium in mixed strategies, which are represented as mixed quantum states, by employing Glicksberg’s [58] extension of Kakutani’s fixed-point theorem to topological vector spaces. In 2019, Khan et al. found that the Kakutani fixed-point theorem does not directly apply to quantum games involving pure quantum strategies [59]. However, by using Nash’s embedding theorem, which embeds compact Riemannian manifolds into Euclidean space [60], and under certain conditions, the Kakutani fixed-point theorem can be indirectly applied to ensure Nash equilibrium in pure quantum strategies. They did a formal mathematical discussion of non-cooperative game theory and fixed points (see Ref. [59] for details).
@@
Our contribution In our game, several key points should be noted as it is a mixed quantum strategic game. When players can choose their quantum strategies based on a probability distribution, thereby using mixed quantum strategies, Meyer demonstrated through Glicksberg’s fixed-point theorem [58] that a Nash equilibrium will always be present. Meyer’s research also offers insights into the equilibrium behavior of quantum computational mechanisms. His approach to identifying quantum advantage as a Nash equilibrium in quantum games remains largely unexplored [61]. Additionally, in the realm of quantum communication protocols, where quantum processes are typically noisy and represented as density matrices or mixed quantum states, the Meyer–Glicksberg theorem ensures the existence of a Nash equilibrium [11]. Generally, in quantum communication protocols, less restricted quantum players can utilize their mixed strategies to calculate their individual best response functions based on their payoffs. The solutions of best response functions give the Nash equilibrium points. While not all of these points may be Pareto optimal Nash equilibrium points, they serve as a basis for determining the probabilities of mixed strategies employed by the players. This, in turn, allows for an investigation into the existence of Pareto optimal Nash equilibrium points. The outcome of this analysis helps to determine the secure bounds for various quantum information parameters (e.g., secret key rate, QBER) with appropriate use of the fundamental concepts of quantum information theory [62, 63, 64, 65]. In our letter, we take the DL04 protocol [26] and analyze it as a game, considering different types of attack (collective attacks and IR attack) to get the secure threshold limit for QBER.
@@
In the DL04 protocol, two quantum players employ strategies to ensure secure communication. Our focus lies in determining the secure bound on QBER when accounting for the presence of an eavesdropper (Eve), who may utilize both quantum and classical attack strategies. We consider that the quantum Eve employs the following attack strategies: Wójcik’s original attack [73], Wójcik’s symmetrized attack [73] and Pavičić attack [74], denoted as $E_{1}$, $E_{2}$ and $E_{3}$, respectively. Additionally, there is a classical Eve employing an Intercept Resend (IR) attack strategy ($E_{4}$). The detailed analysis of attack strategies is conducted in Appendices A, B, C and D. Here, we briefly describe Eve’s attack strategies to aid readers in understanding the concepts (for more details, see the Appendices). In the $E_{1}$ attack, Eve utilizes the unitary operation $Q_{txy}$ during the B-A attack and its conjugate, $Q_{txy}^{\dagger}\left(\equiv Q_{txy}^{-1}\right)$, during the A-B attack, where $Q_{txy}(={\rm SWAP}_{tx}\,{\rm CPBS}_{txy}\,H_{y}\equiv{\rm SWAP}_{tx}\otimes I_{y}\,{\rm CPBS}_{txy}\,I_{t}\otimes I_{x}\otimes H_{y})$. Here, A-B and B-A attack refer to Eve’s attack on the quantum channel when state transmission is from Alice to Bob and from Bob to Alice, respectively. This attack results in the state being in a higher dimensional Hilbert state with Eve’s unitary operations, leading to a higher degree of randomization through quantum superposition. This affects the final joint probabilities of Alice’s, Bob’s and Eve’s measurement outcomes as $p_{000}^{E_{1}}=q,$ $p_{001}^{E_{1}}=p_{010}^{E_{1}}=p_{011}^{E_{1}}=0,$ $p_{100}^{E_{1}}=\frac{1}{4}\left(1-q\right),$ $p_{101}^{E_{1}}=\left(1-q\right)\left(\frac{1}{4}+\frac{p}{2}\right),$ $p_{110}^{E_{1}}=0$ and $p_{111}^{E_{1}}=\frac{1}{2}\left(1-p\right)\left(1-q\right)$. The QBER in the $E_{1}$ attack is $\frac{1}{2}\left(1-q\right)\left(1+p\right)$ and the detection probability of Eve’s presence is 0.1875. The $E_{2}$ attack is similar to the $E_{1}$ attack, with the only difference being that with a probability of $\frac{1}{2}$, an additional unitary operation $S_{ty}$ is applied right after the operation $Q_{txy}^{-1}$ during the A-B attack. The $S_{ty}$ operation is defined as $X_{t}Z_{t}{\rm CNOT}_{ty}X_{t}$. In this scenario, the final joint probabilities of Alice’s, Bob’s and Eve’s measurement outcomes are $p_{000}^{E_{2}}=\frac{1}{2}\left[\frac{q}{4}\left(1+p\right)+q\right]=\frac{q}{8}\left(5+p\right),$ $p_{001}^{E_{2}}=\frac{q}{8}\left(1+p\right),$ $p_{010}^{E_{2}}=p_{011}^{E_{2}}=\frac{q}{8}\left(1-p\right),$ $p_{100}^{E_{2}}=\frac{1}{4}\left(1-q\right),$ $p_{101}^{E_{2}}=\frac{1}{4}\left(1-q\right)\left(1+2p\right),$ $p_{110}^{E_{2}}=0$ and $p_{111}^{E_{2}}=\frac{1}{2}\left(1-p\right)\left(1-q\right)$. The QBER in the $E_{2}$ attack is $\frac{1}{4}\left(2+2p-q-3pq\right)$ and the detection probability of Eve’s presence is 0.1875. In the $E_{3}$ attack, similar to the previous attacks, Eve applies the unitary operation $Q_{txy}^{\prime}$ when the photon is traveling from Bob to Alice in the B-A attack scenario. Conversely, in the A-B attack scenario, Eve applies the inverse of $Q_{txy}^{\prime}$ $\left(Q_{txy}^{\prime-1}\right)$ to the traveling photon ($t$), where $Q_{txy}^{\prime}={\rm CNOT}_{ty}\left({\rm CNOT}_{tx}\otimes I_{y}\right)\left(I_{t}\otimes{\rm PBS}_{xy}\right){\rm CNOT}_{ty}\left({\rm CNOT}_{tx}\otimes I_{y}\right)\left(I_{t}\otimes H_{x}\otimes H_{y}\right)$. In this scenario, the final joint probabilities of Alice’s, Bob’s and Eve’s measurement outcomes are $p_{000}^{E_{3}}=q,$ $p_{001}^{E_{3}}=p_{010}^{E_{3}}=p_{011}^{E_{3}}=0,$ $p_{100}^{E_{3}}=p_{101}^{E_{3}}=p_{110}^{E_{3}}=0,$ and $p_{111}^{E_{3}}=\left(1-q\right)$. The QBER in the $E_{3}$ attack is 0 and the detection probability of Eve’s presence is 0.1875. The $E_{4}$ attack is essentially an IR attack, where the final joint probabilities of Alice’s, Bob’s, and Eve’s measurement outcomes are $p_{000}^{E_{4}}=\frac{3}{4}q,$ $p_{001}^{E_{4}}=p_{011}^{E_{4}}=0,$ $p_{010}^{E_{4}}=\frac{1}{4}q,$ $p_{100}^{E_{4}}=p_{110}^{E_{4}}=0,$ $p_{101}^{E_{4}}=\frac{1}{4}\left(1-q\right)$ and $p_{111}^{E_{4}}=\frac{3}{4}\left(1-q\right)$. The QBER in the $E_{4}$ attack is 0.25 and the detection probability of Eve’s presence is 0.375.
Before describing the conventional payoff function, it is crucial to map the DL04 protocol in a game-theoretic scenario where principles of game theory is applied into a quantum communication protocol (see Figure 1). For simplicity, we divide DL04 quantum game into different game scenarios. However, in what follows, we explain the mapping in a general context [55, 12, 75]. Let us consider game $\mathcal{G}$ as a normal form game. This game is defined as a function with an appropriate domain and range, with players’ preferences defined over the elements of the range, inducing rational strategic choices within the domain. We now describe the players’ choices in the quantum domain. In more general terms, the game $\mathcal{G}$ has a range that is equal to the output set $O$ and a domain that is the Cartesian product $S_{1}\times S_{2}\times\cdots\times S_{n}$, where $S_{i}$ represents the set of pure strategies for the $i^{th}$ player. Each element of the domain set $S_{1}\times S_{2}\times\cdots\times S_{n}$ can be referred to as a play of game or a strategy profile. Formally, a game $\mathcal{G}$ is a function, $\mathcal{G}:S_{1}\times S_{2}\times\cdots\times S_{n}\longrightarrow O$. In this context, $n$ represents the number of players. Since our game involves three participants, Alice, Bob and Eve. Bob’s choice of prepared state, Alice’s choice of measurement basis and Eve’s choice of attack strategy are within the domain set $S_{1}\times S_{2}\times S_{3}$, where $S_{1},S_{2}$ and $S_{3}$ are the strategy choices of Alice, Bob and Eve, respectively. Thus, the function $\mathcal{G}$ can be specifically expressed as $\mathcal{G}:S_{1}\times S_{2}\times S_{3}\longrightarrow O$ in the form of pure strategies. Here, Bob prepares quantum states in the $Z$ or $X$ basis and sends them to Alice. For a security check, Alice measures part of Bob’s sequence and then encodes a message by performing the Pauli operation before sending it back to Bob. During this process, Eve employs various attack strategies through the quantum channel. In what follows, we elucidate how this game is incorporated into the quantum communication protocol.
Alice arranges her initial sequence as a mixed state, represented by $\frac{p}{2}\left(|0\rangle\langle 0|+|1\rangle\langle 1|\right)+\frac{1-p}{2}\left(|+\rangle\langle+|+|-\rangle\langle-|\right)$. The states of Alice resides within a 2-dimensional projective complex Hilbert space, denoted as $\mathcal{H}_{2}$. This mixed state contains with four pure states, which can be interpreted as a superposition state within $\mathcal{H}_{2}$. This implies that if $|0\rangle$ undergoes a projective measurement in the $X$ basis [^1] Let us assume that the process of quantum measurement in a 2-dimensional space can be characterized by the operator set {$M_{0}$,$M_{1}$: $M_{0}=|+\rangle\langle+|$, $M_{1}=|-\rangle\langle-|$}., it will collapse into |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ with a probability of $\frac{1}{2}$. This occurrence is a result of quantum superposition. Eve also executes attacks using quantum unitary gates and ancilla states, which operate within the projective complex Hilbert space. Specifically, the Eve’s $E_{1}$, $E_{2}$ and $E_{3}$ attacks operate within an 8-dimensional Hilbert space $\mathcal{H}_{8}$, and the $E_{4}$ attack operates within $\mathcal{H}_{2}$ space. In our different game scenarios, we assume that Eve attacks with $E_{i}$ and $E_{j}$ with probabilities $r$ and $1-r$, respectively. As a result of these operations by Eve, the states of Alice and Bob are further increase the quantum superposition based on the unitary operators utilized in Eve’s attacks ( inplying that the basis used by alice total number of possible outcome increases). This increase of superposition impacts the measurement outcome of the legitimate player in this game scenario [^2] The detailed descriptions of Alice’s and Bob’s unitary operations and the results of their measurements are elaborately presented in the appendices. that also increase the range of the game $\left(O\right)$, for $O=\left\{o_{1},o_{2},o_{3}\right\}$, $o_{1},o_{2}$ and $o_{3}$ are the output of Alice, Bob and Eve, respectively. Alice employs Pauli operations to encode the message, which also exists within $\mathcal{H}_{2}$ space. Consequently, we can describe our game as a mixed strategy game, defined as a function $\mathcal{M}$. The 1--simplex $\Delta_{1}$, represents the set of probability distributions over each player’s pure strategies, also known as the set of mixed strategies. Thus, our mixed strategy game $\mathcal{M}$ is defined with a domain equal to the Cartesian product of the sets of probability distributions over the pure strategies of the players. Formally, this is represented as:
$\begin{array}[]{lcl}\mathcal{M}&:&\Delta_{1}\times\Delta_{1}\times\Delta_{1}\longrightarrow\Delta_{7}\\ \\ \mathcal{M}&:&\left(\left(p,1-p\right),\left(q,1-q\right),\left(r,1-r\right)\right)\\ &\mapsto&\left(pqr,pq\left(1-r\right),p\left(1-q\right)r,p\left(1-q\right)\left(1-r\right),\left(1-p\right)qr,\right.\\ &&\left.\left(1-p\right)q\left(1-r\right),\left(1-p\right)\left(1-q\right)r,\left(1-p\right)\left(1-q\right)\left(1-r\right)\right)\end{array},$
Its range includes probability distributions over the outcomes by associating these outcomes with the vertices of the 7--simplex $\Delta_{7}$. The required projective complex Hilbert space for $\mathcal{M}$ is as follows:
$\begin{array}[]{lcl}{\rm Bob^{\prime}s\,prepared\,state}&\longrightarrow&\mathcal{H}_{2}\\ {\rm Alice^{\prime}s\,measuremet\,basis}&\longrightarrow&\mathcal{H}_{2}\\ {\rm Alice^{\prime}s\,unitary\,operation}&\longrightarrow&\mathcal{H}_{2}\\ {\rm Eve^{\prime}s\,}E_{1},E_{2},E_{3}\,{\rm attack}&\longrightarrow&\mathcal{H}_{8}\\ {\rm Eve^{\prime}s\,}E_{4}\,{\rm attack}&\longrightarrow&\mathcal{H}_{2}\end{array}.$
Additionally, we observe that the superposition states belong to Bob, along with the unitary operations executed by Alice and Eve, thus the state belongs within the higher dimensional complex Hilbert space ($\mathcal{H}_{8}$). This process incorporates a higher degree of randomization via quantum superposition. The impact of such high-level randomization reflects on the outputs of the legitimate players or increase the range of this game. Accordingly, these output results influence on the key factors of the players’ payoff function, such as mutual information, QBER and the detection of Eve’s presence. Ultimately, these effects on the payoff functions determine the Nash equilibrium points for the different game scenarios. By employing quantum superposition states and quantum unitary operations, we apply non-cooperative game theory to the DL04 protocol and demonstrate that Nash equilibrium serves as a viable solution concept. Hence, we integrate gaming the quantum in our work [75, 61].
Figure 1: The circuit representation of DL04 quantum game.
@@
We will now define the payoff for each party involved. The mutual information between Alice and Bob has a positive impact on the payoffs of legitimate players (Alice and Bob), but it negatively affects Eavesdropper’s (Eve) payoff. Conversely, the mutual information shared between Alice and Eve, as well as Bob and Eve, negatively affects the payoffs of legitimate players. Additionally, legitimate players benefit from the ability to detect Eve’s presence [^3] The contribution of both parameters is equally significant, contingent upon the specific quantitative values of QBER and the detection probability of Eve’s presence ($P_{d}$) observed in both message and control modes of the DL04 protocol., which improves their payoffs in this competitive scenario. On the other hand, Eve’s payoff increases with the acquisition of more information from Alice and Bob but decreases as the mutual information between Alice and Bob increases. Furthermore, Eve incurs penalties if she is detected. Consequently, Eve’s payoff is enhanced when the probability of remaining undetected is higher. Additionally, Eve may employ various quantum gates to collect information from Alice and Bob, and the greater the number of gates she uses, the more it impacts her overhead, which, in turn, adversely affects her payoff. As a result, the payoff can be customized to reflect the various scenarios and benefits of the players that we intend to analyze. The DL04 protocol, resembling a game, can be conceptualized as a zero-sum game. We may formulate the payoffs of Alice, Bob and Eve [53] for a general attack strategy $\mathcal{E}$ as follows [^4] We assume that the difficulty of preparing quantum states and performing measurement operations is the same for Alice, Bob and Eve. Therefore, we do not factor in this contribution when estimating the payoffs for these parties.,
@@
where $\omega_{a},\,\omega_{b},\,\omega_{c},\,\omega_{d},\,\omega_{e},\,\omega_{f},\,\omega_{g},\,\omega_{h},\,\omega_{i},\,\omega_{j}$ and $\omega_{k}$ are all positive real numbers, and they are interpreted as the factors or weights assigned to each component within the payoff such that $\stackrel{{\scriptstyle[}}{{m}}=a]{d}{\sum}\omega_{m}=1$ and $\stackrel{{\scriptstyle[}}{{n}}=e]{k}{\sum}\omega_{n}=1$ where $m\in\{a,b,c,d\}$ and $n\in\{e,f,\ldots,k\}$ Further, $I\left({\rm A,B}\right)$, $I\left({\rm A,E}\right)$ and $I\left({\rm B},{\rm E}\right)$ are the mutual information between Alice and Bob, Alice and Eve and Bob and Eve, respectively. Here, $P_{d}$ is the probability of detection of Eve’s presence, $n_{1}$, $n_{2}$ and $n_{3}$ are the number of single-qubit, two-qubit and three-qubit gates, respectively. Here, it may be noted that in the payoff function described above, $P_{d}$ and $QBER$ are considered on equal footing because both $P_{d}$ and $QBER$ reveal the presence of Eve with a similar effect.
Assuming Eve has unlimited quantum resources and perfect quantum gates, we can set $\omega_{i}=\omega_{j}=\omega_{k}=0$. This implies she can use as many quantum gates as needed for her attack without negatively impacting her payoff function. To simplify the analysis, in what follows we have considered the effect of each component in the payoff is the same. Thus, in our consideration, $\omega_{a}=\omega_{b}=\omega_{c}=\omega_{d}=\omega_{e}=\omega_{f}=\omega_{g}=\omega_{h}=0.25$. The Eq. (1) is modified as,